OpenStudy (twistrose):

Solve for x: -2/3(3x-4)+3x=5/6

Awolflover1 (awolflover1):

\[\frac{ -2 }{ 3 } (3x−4)+3x= \frac{ 5 }{ 6 }\]

OpenStudy (opticalmango81):

I am not good at math lol but i tried to think but all im getting are brainfarts lol

Awolflover1 (awolflover1):

\[x+\frac{ 8 }{ 3 }= \frac{ 5 }{ 6}\]

Awolflover1 (awolflover1):

Subtract \[\frac{ 8 }{ 3 }\] from both sides.

OpenStudy (twistrose):

is it -11/6?

Awolflover1 (awolflover1):

correct (:

OpenStudy (twistrose):

ok thanks